OpenStudy (anonymous):

find the equation of the line that pasess through the pair of points (3,7) and (-6,-2) in standard form

OpenStudy (anonymous):

First find the slope

OpenStudy (anonymous):

\[m=\frac{ y2-y1 }{ x2-x1 }\] \[y-y1=m(x-x1)\]

OpenStudy (anonymous):

Then use point slope form as surdawi suggested

OpenStudy (anonymous):

thanx guys

OpenStudy (anonymous):

y=x+4